Please Solve by Elimination
-3x-5y+3z=-2
4x+4y+2z=6
x-2y-3z=-5
Please help me to figure out this problem, I dont even know where to begin.

Answer by ankor@dixie-net.com(22740)

Solve by Elimination
-3x - 5y + 3z =-2
4x + 4y + 2z = 6
x - 2y - 3z =-5
:
Multiply the 3rd equations and add
-3x - 5y + 3z =-2
4x + 4y + 2z = 6
-x + 2y + 3z = 5
---------------------adding eliminates x
0 + y + 8z = 9
y = (9-8z)
:
Multiply the 3rd equation by 3 add to the first equation
-3x - 5y + 3z =-2
3x - 6y - 9z =-15
---------------------adding eliminates x
0 - 11y - 6z = -17
replace y with (9-8z)
-11(9-8z) - 6z = -17
-99 + 88z - 6z = -17
82z = -17 + 99
82z = 82
z = 1
:
find y:
y = (9-8z)
y = 9 - 8(1)
y = 1
:
I'll let you find x, check the solutions in the original equations
(Don't be surprised if x=0)